OpenStudy (anonymous):

You are going to design an advertisement for a new polynomial identity that you are going to invent. Your goal for this activity is to demonstrate the proof of your polynomial identity through an algebraic proof and a numerical proof in an engaging way! Make it so the whole world wants to purchase your polynomial identity and can't imagine living without it! You may do this by making a flier, a newspaper or magazine advertisement, making an infomercial video or audio recording, or designing a visual presentation for investors through a flowchart or PowerPoint.

jimthompson5910 (jim_thompson5910):

First thing you need is a polynomial identity.

jimthompson5910 (jim_thompson5910):

Do you have one in mind?

OpenStudy (anonymous):

No I dont

jimthompson5910 (jim_thompson5910):

do you know what an identity is?

OpenStudy (anonymous):

..no

jimthompson5910 (jim_thompson5910):

an identity is an equation that is always true for all allowed values in the domain

jimthompson5910 (jim_thompson5910):

something like 1 = 1, 10 = 10, x = x are all identities

OpenStudy (anonymous):

So I need to create one?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

another identity: x+x = 2x another one: 23x - 8x = 15x

jimthompson5910 (jim_thompson5910):

all of the equations posted are true for all values of x

OpenStudy (anonymous):

How about 13x+4x= 17x?

jimthompson5910 (jim_thompson5910):

that's an identity, good

jimthompson5910 (jim_thompson5910):

whatever you plug in for x, it will be a true equation

OpenStudy (anonymous):

Alright what do I do next?

jimthompson5910 (jim_thompson5910):

something more complicated x^2 + 3x - 2x + 9 = x^2 + x + 9

jimthompson5910 (jim_thompson5910):

well the algebraic proof for 13x+4x= 17x is simply showing that both sides are equal to the same thing

jimthompson5910 (jim_thompson5910):

numerically, you just plug in various numbers to show that the equation is always true

OpenStudy (anonymous):

So now I would create a more complex one?

jimthompson5910 (jim_thompson5910):

you can create one as simple or as complex as you want

jimthompson5910 (jim_thompson5910):

all they care about is that you have an identity of some sort (that simplifies to a polynomial)

OpenStudy (anonymous):

So is the one I did fine ? Or would it be better if I did a more complex one?

jimthompson5910 (jim_thompson5910):

it might be better to go with a complicated one since you have more work to show

jimthompson5910 (jim_thompson5910):

Then you have to design some sort of flier or presentation to show the world this identity and you have to make it like a sales pitch it seems

OpenStudy (anonymous):

I'm REALLY bad at math so could you help me make it step by step?

jimthompson5910 (jim_thompson5910):

which identity are you going with?

OpenStudy (anonymous):

(ax+b)

jimthompson5910 (jim_thompson5910):

remember, an identity is an equation that's always true (for all of the variables' values)

OpenStudy (anonymous):

oh so I'll use the one I made earlier

jimthompson5910 (jim_thompson5910):

here's a good somewhat complicated identity

jimthompson5910 (jim_thompson5910):

(x+3)^2 = x^2 + 6x + 9

jimthompson5910 (jim_thompson5910):

no matter what you plug in for x, you'll get the same number on both sides (when you evaluate), so you'll get a true equation for all real numbers x

OpenStudy (anonymous):

Alright so we got that, now what?

jimthompson5910 (jim_thompson5910):

you have to prove that (x+3)^2 = x^2 + 6x + 9 is true

jimthompson5910 (jim_thompson5910):

both algebraically and numerically

OpenStudy (anonymous):

So pick a number to be x?

jimthompson5910 (jim_thompson5910):

for the numeric proof, yes

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

ok replace every x with 4 and see if you get a true equation

OpenStudy (anonymous):

4+3)^2 = 4^2 + 6(4) + 9

jimthompson5910 (jim_thompson5910):

(4+3)^2 = 4^2 + 6(4) + 9 good

jimthompson5910 (jim_thompson5910):

then evaluate each side

OpenStudy (anonymous):

what does ^ stand for ?

jimthompson5910 (jim_thompson5910):

^ means exponent

jimthompson5910 (jim_thompson5910):

so 8^2 means 8 exponent 2 or 8 squared

jimthompson5910 (jim_thompson5910):

\(\Large 8\wedge 2 = 8^2\)

OpenStudy (anonymous):

ohhh okay

jimthompson5910 (jim_thompson5910):

It's mainly notation for typing in text only (like in a text document or on a forum like this) and calculators also use this notation too

OpenStudy (anonymous):

49 is the first half?

jimthompson5910 (jim_thompson5910):

left side, yes

OpenStudy (anonymous):

I got 41 for the other side

OpenStudy (anonymous):

oh no i did it wrong hold on

OpenStudy (anonymous):

I got 49 got the other side

jimthompson5910 (jim_thompson5910):

you should get 49 for both sides, so be extra sure you get 49 for the RHS (right hand side)

OpenStudy (anonymous):

I did :)

OpenStudy (anonymous):

What now?

jimthompson5910 (jim_thompson5910):

use other numbers to show it works for them as well I'd say 3 examples is probably good enough but you can go as high as you want

jimthompson5910 (jim_thompson5910):

this will all be for the numeric form of the proof

OpenStudy (anonymous):

Alright I'll do that now.

OpenStudy (anonymous):

7+3)^2 = 7^2 + 6(7) + 9

OpenStudy (anonymous):

100 on the first side

OpenStudy (anonymous):

I got 100 for both sides

jimthompson5910 (jim_thompson5910):

(7+3)^2 not 7+3)^2

jimthompson5910 (jim_thompson5910):

good, 100 = 100 is the final result, so that verifies the identity for x = 7

OpenStudy (anonymous):

whoops

jimthompson5910 (jim_thompson5910):

Here is one way to algebraically prove it's an identity

jimthompson5910 (jim_thompson5910):

(x+3)^2 = x^2 + 6x + 9 (x+3)(x+3) = x^2 + 6x + 9 x(x+3)+3(x+3) = x^2 + 6x + 9 x^2+3x+3x+9 = x^2 + 6x + 9 x^2 + 6x + 9 = x^2 + 6x + 9

jimthompson5910 (jim_thompson5910):

so that takes care of the algebraic approach

OpenStudy (anonymous):

Alright, thank you !

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Thats everything right ?

jimthompson5910 (jim_thompson5910):

you just have to make the presentation, but that's it in terms of the math portion really

OpenStudy (anonymous):

Alright thank you ! Will you be on here tomorrow?

jimthompson5910 (jim_thompson5910):

I should be, so feel free to message whenever

OpenStudy (anonymous):

I will, again thank you sooo much !

jimthompson5910 (jim_thompson5910):

sure thing